The supplement of an angle added to the complement of the angle gives 210 degrees. What is the measure of the angles?
Answer by TimothyLamb(4379)

x = angle measure in degrees
---
(180 - x) + (90 - x) = 210
270 - 2x = 210
2x = 270 - 210
x = (270 - 210)/2
(270 - 210)/2
---
copy and paste the above expression in to this calculator:
https://sooeet.com/math/scientific-calculator.php
---
x = 30
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---